trying to understand how to do this? my teacher isnt giving us notes and i dont know how to do it.
give 3 solutionsfor the following equations.
y = x + -8

Answer by stanbon(75887)

trying to understand how to do this? my teacher isnt giving us notes and i dont know how to do it.
give 3 solutionsfor the following equations.
y = x + -8
----
x is called "the independent variable".
So you can give any value you want to to "x".
Then you can solve for "y".
-----
If x = 3, y = 3-8 = -5, giving you pair (3,-5)
If x = -2, y = -2-8 = -10, giving you pair (-2,-10)
If x = 0, y = 0-8 = -8, giving you pair (0,-8)
=======================
Cheers,
Stan H.
==========